# FlightScope: A Deep Comprehensive Review of Aircraft Detection Algorithms in Satellite Imagery

### v3

@@ 1 Introduction
Remote sensing plays a fundamental role in acquiring information about the Earth’s surface using various types of vision sensors such as Thermal Infra-Red (TIR) and RGB cameras [1]. This field encompasses a wide range of technologies and methodologies aimed at capturing, analyzing, and interpreting data from various sources. Within the realm of remote sensing, one of the most significant applications is the detection and localization of small objects [2]. Object detection from satellite imagery holds great importance in various domains, such as defence and military applications [3], urban studies [4], airport surveillance, vessel traffic monitoring [5] and transportation infrastructure determination [6, 7]. Unlike photographic pictures, remote sensing images obtained from satellite sensors are more difficult to interpret due to factors such as atmospheric interference, viewpoint variation, background clutter, and illumination differences [8, 9]. Additionally, satellite images cover larger areas, typically surface like 10 $\times$ 10 km2 per frame, representing the intricate landscape of the Earth’s surface with two-dimensional images that possess less spatial detail compared to digital photographs from cameras.
@@
Traditional approaches to aircraft detection relied on manual feature engineering and machine learning techniques. However, these methods often struggles to handle the complexities of satellite imagery and achieve high accuracy. The advent of deep learning, particularly Convolutional Neural Networks (CNNs), has revolutionized object detection tasks by enabling the automatic extraction of intricate visual representations. One notable deep learning method for object detection is the You Only Look Once (YOLO) [10]. YOLO divides the input image into a grid and predicts bounding boxes and class probabilities directly from the grid cells. Its architecture allows for real-time detection and can handle multiple object classes simultaneously. However, YOLO may encounter challenges in accurately localizing small objects due to the grid cell structure and limited receptive fields. Another popular deep learning technique is the Single Shot MultiBox Detection (SSD) approach [11], which utilizes a series of convolutional layers to generate a diverse set of default bounding boxes at different scales and aspect ratios. By applying a set of predefined anchor boxes to the feature maps, SSD performs multi-scale object detection efficiently. However, it may face difficulties in detecting small objects and suffers from a large number of default boxes, leading to computational overhead. Region-based CNNs (RCNN) [12] and its variants, such as Fast RCNN [13], Faster RCNN [14] and Mask RCNN [15], have also been widely used for object detection tasks. These methods employ a two-stage approach, in which potential object regions are first proposed and then classified and refined. By utilizing region proposals, these methods achieve accurate localization and exhibit strong performance. However, they are computationally expensive and slower than one-stage detectors such as YOLO and SSD. In addition to YOLO, SSD, and RCNN variants, there are several other deep learning methods that have been explored for object detection, such as EfficientDet [16], RetinaNet [17], and CenterNet [18]. More recently, the success of attention mechanism in the field of language processing has received a lot of consideration and has been brought into the field of computer vision as well [19]. This rise has led to an interesting improvement of performance in many computer vision subfields, including image classification [20] and object detection [21]. Each of these methods has its own unique architecture and advantages, aiming to improve both accuracy and efficiency in object detection tasks.
@@
Models Tested Short Description
YOLO v5 [22] Real-time object detection model that processes images in one pass through a CNN, widely used for various applications including autonomous vehicles, surveillance systems, and robotics.
@@
YOLO v8 [10] Updated version of the YOLO object detection system, incorporating advancements in network architecture and training techniques to achieve better efficiency in real-time object detection tasks.
@@
RTMDet [23] An improved CNN network for higher accuracy while maintaining the same real-time performance of the YOLO model series.
@@
DETR [24] New object detection architecture based on transformers and attention mechanisms. It has proved its efficiency on the COCO dataset.
@@
Faster RCNN [14] Region-Based CNN is an architecture for object detection that has proven its efficiency in accurate bounding box extraction. Faster RCNN is an algorithm that improves the detection speed of RCNNs.
@@
SSD [11] Another real-time object detection model that predicts bounding boxes from features map at multiple scales.
@@
CenterNet [18] This model, initially based on the CornerNet [25] architecture, has been able to achieve state-of-the-art performance on COCO dataset for object localization in term of real-time and accuracy.
@@
RetinaNet [26] Real-time object detection model that addresses the imbalance between foreground and background examples during training using a novel focal loss function.
@@
The objective of this study is to benchmark and compare multiple state-of-the-art object detection methods prepared and trained specifically for the use case of aircraft detection in satellite images. The focus is on their application in satellite surveillance, and air traffic management. This paper significantly contributes to the research on satellite imagery analysis by implementing, training and validating one variant of the eight leading object detection neural network architectures listed in Table 1, alongside other cutting-edge deep learning architectures. The authors have directly implemented and tested each network to control the testing conditions and prevent bias or spurious outcomes that might arise. In addition to a detailed overview of object detection models that is fundamental for both researchers and practitioners, this work provides also a thorough examination of their performance, precision, and computational complexity. The findings provide critical information that improves the selection process for the most efficient aircraft detection methods in satellite imagery. This is supported by comprehensive training and initial validation on the HRPlanesv2 and further validation on GDIT datasets (for details on the datasets see Section 3), marking an important advancement in the study of the precision and efficiency of remote sensing technologies and their application in real-world scenarios, thus underlining its substantial relevance and potential impact on future research in the field.
@@
Additionally, to make a more useful contribution to the computer vision community, all the code used for the benchmarks in this paper is available, and could be reproduced from the GitHub repository at https://github.com/toelt-llc/FlightScope_Bench.
@@
Following the introduction, the rest of the paper is structured as follow: similar comparative studies are discussed and reviewed in Section 2. An overview of existing datasets including airplanes in satellite images is discussed in Section 3, while in Section 4 a more detailed state-of-the-art description of the object detection architectures (listed in Table 1) is given. Section 5 present the setup of the benchmark, along with the results and discussion. Finally, Section 6 concludes and summarizes the findings of this comparative study.
@@ 2 Related work
Many remote sensing methods and models have been studied and proposed during the past decade in various fields such as environmental monitoring [27], object and image Geo-localization [28, 29], urban planning [30], and agriculture [29]. In the context of object detection, the methods that are usually proposed are trained, validated and tested on images gathered from ground and usual vision sensors. On those types of images, good performances are observed, therefore, the need for performance evaluation of remote sensing types of images. Remote sensing imagery typically encompasses vast areas with varying resolutions, making the detection of small objects, such as vehicles or infrastructure, particularly challenging [31]. Additionally, factors such as varying illumination conditions, occlusions due to weather, and object scale variations further complicate the task.
@@
A similar study to this work, conducted by Alganci, et. al. [9] delves into the detection of small objects from satellite imagery. This research focusses on evaluating the performance of three state-of-the-art convolutional neural network (CNN)-based object detection models specifically tailored for identifying airplanes in very high-resolution (VHR) satellite images. The authors underscore the importance of accurate and efficient detection methods in satellite imagery due to its large data size and expansive aerial coverage. For their study, the authors have used the DOTA dataset [32], a multiple classes open-source repository explicitly created for object detection in remote sensing images. This dataset encompasses satellite image patches sourced from platforms such as Google Earth, Jilin 1 (JL-1), and Gaofen 2 (GF-2) satellites, featuring 15 object categories, among which is the airplane class. The comparative evaluation conducted in [9] assesses three object detection models: Faster R-CNN [14], SSD [11], and YOLO-v3 [33], using the DOTA dataset for both training and testing. Performance metrics including COCO metrics, F1 scores, and processing time are employed for evaluation. The summary of this work reveal that Faster R-CNN exhibits superior detection accuracy, with YOLO-v3 showcasing faster convergence capabilities. SSD, although proficient in object localization, faces challenges with training convergence.
@@
Additionally, another study presented in [34] focusses on the development and evaluation of the first version of the HRPlanes benchmark dataset for deep learning-based airplane detection using satellite imagery from Google Earth. The authors describe the HRPlanes dataset and some of the images captured by different satellites to represent diverse landscapes, seasonal variations, and satellite geometry conditions. The dataset is then selected for training and validation of two widely used object detection methods, YOLOv4 [35] and Faster R-CNN [14]. The comparative study between YOLOv4 and Faster R-CNN in the context of airplane detection from satellite imagery reveals interesting findings. The study highlights that the boundaries of bounding boxes for YOLOv4 are better at certain scales compared to Faster R-CNN. For instance, in some cases, YOLOv4 performs better in detecting small airplanes, while Faster R-CNN excels in detecting larger ones. The results also indicate that YOLOv4 is more effective in creating accurate bounding boxes for commercial planes in large-scale imagery, possibly due to the presence of boarding bridges near the planes. Additionally, both deep learning models demonstrate the ability to detect moving planes, even in scenarios with motion blur effects in the images.
@@ 3 Aircraft Datasets
In the field of aircraft detection and remote sensing, access to high-quality and diverse datasets is important for the development and evaluation of computer vision algorithms. In some comparative studies, DOTA dataset [32] is selected, this latter encompasses objects other than aircrafts such as airports, bridges and containers. This section reviews a collection of datasets specifically and exclusively designed for aircraft detection research, each offering unique features and advantages. An overview of these datasets is given in Table 2.
@@ 4 Literature Review: Object Detection
Many object detection models have been proposed in the past decades, the majority are based on CNNs. Researchers have so far classified those models into two primary categories: one-stage and two-stage architectures [42, 43]. However, the rise of self-attention mechanism [44] recently has led to a new category of network architectures that are based on transformers. Fig. 1) group some of the most used and known models along with two types of classifications of deep learning models: (1) either categorized based on their network architecture (one-stage, two-stages, or transformer), or (2) based on their performances in real-time and detection accuracy highlighted in respectively blue and yellow colors. The overlapping area drawn in green feature the object detection models with a balanced performances in both real-time detection and accuracy.
One-stage models [45] have been mainly known for their real-time deployment performances, because they are usually not computationally extensive and a single pass through the network is sufficient to produce estimations of object bounding boxes. However, the main limitation in this category is the detection accuracy, which might not be enough for some applications requiring very high detection confidence. In the case of a two-stage object detection models, an additional stage is introduced to generate generic object proposals which make the model more efficient in its detection operation [46]. The purpose of this stage is to produce candidate bounding boxes, which may not be highly accurate, and effectively exclude background areas from further processing. Subsequently, the next stage of the model undertakes the more computationally intensive tasks of classifying objects and refining the bounding boxes generated by the previous stage. The third category is transformer-based architecture [47], which seems to be a good balance between accuracy on real-time detection as it has been tested and compared on common datasets. This category have been tested on common large datasets and it leverage the use of the self-attention mechanism to produce reliable and accurate bounding box estimations while being able to perform real-time detection.
One-stageDetection AccuracyReal-time inferenceTransformerTwo-stages Overlapping Deep learning based Object Detection CenterNet [18]CornerNet [25]YOLO [10]RetinaNet [26]RTMDet [18]SSD [11]RefineDet [48]Casc.-RCNN [49]R-FCN [50]RCNN [51]Fast-RCNN [13]Faster-RCNN [14]Mask-RCNN [15]DETR [52]RT-DETR [24]ViTDet [53]
Figure 1: Classification of object detection methods based on: (1) their architecture (One-stage, two-stages and transformer network); (2) on detection accuracy (in orange) and real-time detection (blue). The red dot highlights the models that are implemented, trained and validated in this work. The green outline (indicated as ‘Overlapping’ in the image) groups the models that usually perform well in both accuracy and inference time response.
The choice between object detection architecture depends on the application and the nature of images to process which involves a trade-off between speed and accuracy [54]. One-stage models tend to offer faster processing speeds but may exhibit lower accuracy compared to their two-stage counterparts. The advantages of one-stage algorithms could be summarized as follow [55, 56, 57]:
@@
  • Real-time Performance: One-stage detectors are designed to achieve real-time or near real-time performance, making them suitable for applications where fast inference is crucial. These algorithms are commonly used in scenarios that require quick responses, such as autonomous driving, video analysis, and robotics [58].
  • Higher Recall: One-stage detectors tend to have a higher recall rate compared to two-stage detectors [59, 11]. They are capable of detecting a larger number of objects in an image, including small or occluded objects, due to their dense and dense-like prediction strategies. This higher recall can be advantageous in applications where comprehensive object detection is more important than achieving extremely precise bounding box localization.
@@
On the other hand, two-stage models generally achieve higher accuracy but sacrifice some speed due to the additional processing stage. The advantages of two-stage algorithms over the others could be summarized into the following points:
@@
  • Sampling efficiency: Two-stage detectors employ a sampling mechanism to select a sparse set of region proposals, effectively eliminating a significant portion of the negative proposals. Conversely, one-stage detectors take a different approach by directly considering all regions in the image which sometimes introduce class imbalance [60].
@@
  • Feature extraction: Two-stage detectors can allocate a larger head network for proposal classification and regression as they only process a small number of proposals. This allows for the extraction of richer features, contributing to improved performance.
@@
  • Recall: Two-stage detectors leverage the RoIAlign [15, 61] operation to extract high-quality features from each proposal, ensuring location consistency. Conversely, in one-stage detectors, different region proposals may share the same feature, leading to coarse and spatially implicit representations that can cause feature misalignment.
@@
  • Accuracy: Two-stage detectors refine the object locations twice, once in each stage. Consequently, the bounding boxes generated by these models exhibit better accuracy compared to one-stage methods but at the expense of real time performance.
@@
Consequently, the trade-off between one-stage, two-stage and transformer-based architectures necessitates careful consideration based on the specific requirements of the application at hand as well as the nature of images. In this context, special attention is directed towards the detection of airplanes within remote sensing images. The choice of algorithms is guided by their widespread usage, popularity in performances when tested on other types of images, and availability as open-source implementations. This will include YOLO, CenterNet, RTMDet, SSD, RetinaNet, Faster-RCNN and DETR.
4.1 You Only Look Once

### v4

# FlightScope: An Experimental Comparative Review of Aircraft Detection Algorithms in Satellite Imagery

@@ 1 Introduction
Remote sensing plays a fundamental role in acquiring information about the Earth’s surface using various types of vision sensors such as Thermal Infra-Red (TIR), synthetic aperture radar (SAR), inverse SAR (ISAR) [1] and RGB cameras [2, 3]. This field encompasses a wide range of technologies and methodologies that aim to capture, analyze, and interpret data from various sources. Within the realm of remote sensing, one of the most significant applications is the detection and localization of small objects [4]. Object detection from satellite imagery is of great importance in various domains, such as defense and military applications [5], urban studies [6], earth monitoring and assessment [7], airport surveillance and georeferencing [8], vessel traffic monitoring [9] and transportation infrastructure determination [10, 11]. Unlike photographic images, remote sensing images obtained from satellite sensors are more difficult to interpret due to factors such as atmospheric interference, viewpoint variation, background clutter, and illumination differences [12, 13]. Additionally, satellite images cover larger areas, typically 10 $\times$ 10 km2 per frame, representing the intricate landscape of the Earth’s surface with two-dimensional images that possess less spatial detail compared to digital photographs from cameras.
@@
Traditional approaches to aircraft detection relied on manual feature engineering and machine learning techniques. However, these methods often struggle to handle the complexities of satellite imagery and achieve high accuracy. The advent of deep learning, particularly Convolutional Neural Networks (CNNs), has revolutionized object detection tasks by enabling the automatic extraction of intricate visual representations. One notable deep learning method for object detection is You Only Look Once (YOLO) [14]. YOLO divides the input image into a grid and predicts bounding boxes and class probabilities directly from the grid cells. Its architecture allows real-time detection and can handle multiple object classes simultaneously. However, YOLO may encounter challenges in accurately localizing small objects due to the grid cell structure and limited receptive fields. Another popular deep learning technique is Single-Shot MultiBox Detection (SSD) [15], which utilizes a series of convolutional layers to generate a diverse set of default bounding boxes at different scales and aspect ratios. By applying a set of predefined anchor boxes to the feature maps, SSD performs multi-scale object detection efficiently. However, it may face difficulties in detecting small objects and suffers from a large number of default boxes, leading to computational overhead. Region-based CNNs (RCNN) [16] and their variants, such as Fast RCNN [17], Faster RCNN [18], and Mask RCNN [19], have also been widely used for object detection tasks. These methods employ a two-stage approach in which potential object regions are first proposed and then classified and refined. Using region proposals, these methods achieve accurate localization and exhibit strong performance. However, they are computationally expensive and slower than one-stage detectors such as YOLO and SSD. In addition to the YOLO, SSD and RCNN variants, there are several other deep learning methods that have been explored for object detection, such as EfficientDet [20], RetinaNet [21], and CenterNet [22]. More recently, the success of the attention mechanism in the field of language processing has received much consideration and has also been brought into the field of computer vision [23]. This rise has led to an interesting improvement in performance in many subfields of computer vision, including image classification [24] and object detection [25]. Each of these methods has its own unique architecture and advantages, aiming to improve both accuracy and efficiency in object detection tasks.
In particular, the objective of this study is to benchmark and compare multiple state-of-the-art object detection methods prepared and trained specifically for the use case of aircraft detection in satellite images, more specifically optical RGB remotely sensed photos. The main focus is to train advanced predictive models and identify highly relevant models for applications in satellite surveillance and air traffic management. This paper significantly contributes to research on satellite imagery analysis by implementing, training, and validating the ten leading object detection neural network architectures listed in Table 1, along with other cutting-edge deep learning architectures. The authors have directly implemented and tested each network to control the testing conditions and prevent bias or spurious outcomes that might arise. The algorithms have been chosen according to two criteria: (i) reference architectures have been selected such as RCNN, YOLO, and more recently DETR as they constitute reference bases for comparisons; (ii) the algorithm must have demonstrated previous state-of-the-art performances on benchmark datasets such as COCO [26] or ImageNet [27].
@@
Models Tested Category Short Description
YOLO v10 [28] Transformer Leverages Transformer-based modules for improved attention and introduces a novel NMS-free strategy for better efficiency. It combines advanced architectural optimization to reduce computational redundancy while maintaining high detection accuracy.
@@
YOLO v8 [14] One-stage Updated version of the YOLO object detection system, incorporating advancements in network architecture and training techniques to achieve better efficiency in real-time object detection tasks.
YOLO v5 [29] One-stage Real-time object detection model that processes images in one pass through a CNN, widely used for various applications including autonomous vehicles, surveillance systems, and robotics.
@@
RTMDet [30] One-stage An improved CNN network for higher accuracy while maintaining the same real-time performance of the YOLO model series.
@@
DETR [31] Transformer New object detection architecture based on Transformers and attention mechanisms. It has proved its efficiency on the COCO dataset.
@@
Faster RCNN [18] Two-stage Region-based CNN is an architecture for object detection that has proven its efficiency in accurate bounding box extraction. Faster RCNN is an algorithm that improves the detection speed of RCNNs.
@@
SSD [15] One-stage Another real-time object detection model that predicts bounding boxes from feature maps at multiple scales.
@@
CenterNet [22] One-stage This model, initially based on the CornerNet [32] architecture, has been able to achieve state-of-the-art performance on the COCO dataset for object localization in term of real-time and accuracy.
@@
RetinaNet [33] One-stage Real-time object detection model that addresses the imbalance between foreground and background examples during training using a novel focal loss function.
Grounding DINO [29] Transformer Cutting-edge Transformer-based open-world object detection model that blends the capabilities of vision Transformers with language understanding. This model excels specifically at zero-shot object detection, enabling it to identify objects not present in its training set using natural language queries.
@@
In addition to a detailed overview of object detection models that is fundamental for both researchers and practitioners, this work also provides a thorough examination of their performance, precision, and computational complexity. The findings provide critical information that improves the selection process for the most efficient aircraft detection methods in satellite imagery. This is supported by comprehensive training and initial validation on the HRPlanesv2 and further validation on GDIT datasets (for details on the datasets, see Section 3), marking an important advancement in the study of the precision and efficiency of remote sensing technologies and their application in real-world scenarios, thus underlining their substantial relevance and potential impact on future research in the field.
@@
Additionally, to make a more useful contribution to the computer vision community, all the code used for the benchmarks in this paper is available and can be reproduced from the GitHub repository at https://github.com/toelt-llc/FlightScope_Bench (accessed on 23 October 2024).
@@
Following the introduction, the rest of this paper is structured as follows: Similar comparative studies are discussed and reviewed in Section 2. An overview of existing datasets including airplanes in satellite images is discussed in Section 3, while in Section 4 a more detailed state-of-the-art description of object detection architectures (listed in Table 1) is given. Section 5 presents the setup of the benchmark, along with the results and discussion. Finally, Section 6 concludes and summarizes the findings of this comparative study.
@@ 2 Related work
Many remote sensing methods and models have been studied and proposed during the past decade in various fields such as environmental monitoring [34], object and image geolocalization [35, 36], urban planning [37], and agriculture [36]. In the context of object detection, the methods that are usually proposed are trained, validated, and tested on images gathered from ground and usual vision sensors. Object detection and segmentation techniques demonstrate accurate predictive performance in these types of images. Consequently, there is a significant need to evaluate the performance of remote sensing images using these methods. Remote sensing imagery typically encompasses vast areas with varying resolutions, making the detection of small objects, such as vehicles or infrastructure, particularly challenging [38]. In addition, factors such as varying illumination conditions, occlusions due to weather, and object scale variations further complicate the task.
@@
A few surveys on small object detection from aerial optical imagery (summary in Table 2) have been proposed, such as ship detection [39], where the authors conduct a critical review of deep learning networks and available datasets. Another paper by W. Chen et al. [40] compared multiple families of RCNN, YOLO, SSD, CornerNet, and other models on their own dataset HRSC2016MS and demonstrated the generalization effectiveness of their model MSSDet based on a ResNet backbone [41]. Ships are relatively easier to detect because of the background aspect of those objects, as in most cases they are found on water, which can be distinguished accurately based on the colors and texture uniformity. A more similar study to this work, conducted by Alganci et al. [13], delves into the detection of small objects from satellite imagery. This research focuses on evaluating the performance of three state-of-the-art Convolutional Neural Network (CNN)-based object detection models specifically tailored for identifying airplanes in very-high-resolution (VHR) satellite images. The authors underscore the importance of accurate and efficient detection methods in satellite imagery due to its large data size and expansive aerial coverage. In another study, the authors used the DOTA dataset [42], a multiple-class open-source repository explicitly created for object detection in remote sensing images. This dataset encompasses satellite image patches sourced from platforms such as Google Earth, Jilin 1 (JL-1), and Gaofen 2 (GF-2) satellites, featuring 15 object categories, among which is the airplane class. The comparative evaluation conducted in [13] assesses three object detection models, Faster R-CNN [18], SSD [15], and YOLO-v3 [43], using the DOTA dataset for training and testing. Performance metrics, including COCO metrics, F1 scores, and processing time, are used for evaluation. The summary of this work reveals that Faster R-CNN exhibits superior detection accuracy, with YOLO-v3 showcasing faster convergence capabilities. SSD, although proficient in object localization, faces challenges with training convergence.
@@
Furthermore, another study presented in [44] focuses on the development and evaluation of the first version of the HRPlanes benchmark dataset for deep learning-based airplane detection using satellite imagery from Google Earth. The authors describe the HRPlanes dataset and some of the images captured by different satellites to represent diverse landscapes, seasonal variations, and satellite geometry conditions. The dataset is then selected for the training and validation of two widely used object detection methods, YOLOv4 [45] and Faster R-CNN [18]. The comparative study between YOLOv4 and Faster R-CNN in the context of airplane detection from satellite imagery reveals interesting findings. The study highlights that the boundaries of bounding boxes for YOLOv4 are better at certain scales compared to Faster R-CNN. For example, in some cases, YOLOv4 performs better in detecting small airplanes, while Faster R-CNN excels in detecting larger ones. The results also indicate that YOLOv4 is more effective in creating accurate bounding boxes for commercial planes in large-scale imagery, possibly due to the presence of boarding bridges near the planes. Additionally, both deep learning models demonstrate the ability to detect moving planes, even in scenarios with motion blur effects in the images.
Table 2: Summary of some reviews of optical remote sensing object detection and classification.
Study Description Summary Dataset Classes
Vehicle and Vessel Detection [46] Investigates the effectiveness of single-shot object detection networks for identifying small objects like vehicles and vessels in satellite imagery with resolutions of 0.3–0.5 m. Key challenges addressed include small object sizes (ranging from 5 pixels to several hundred) and diverse object dimensions. The study annotated 1500 km2 of imagery, equally split between land and water areas. Four models were compared: YOLOv2, YOLOv3, D-YOLO, and YOLT. Extensive hyperparameter tuning was conducted to maximize performance. D-YOLO emerged as the top performer, achieving an average precision (AP) of 60% for vehicles and 66% for vessels. YOLOv3 provided balanced results across metrics, while other models showed varying strengths in detection speed and accuracy. DOTA, OpenSARShip Vehicles, Vessels
Multi-Scale Ship Detection [40] Introduces a framework specifically designed to address the challenges of detecting ships at varying scales in optical remote sensing imagery. Traditional detection methods often struggle with the complexities posed by scale variations, cluttered backgrounds, and low contrast between ships and their surroundings. MSSDet addresses these issues by incorporating a Joint Recursive Feature Pyramid (JRFP) into its architecture, which enhances feature fusion across multiple scales. MSSDet demonstrates robust performance when tested on HRSC2016-MS, the original HRSC2016, and the DIOR dataset, achieving notable mAP improvements over other models like the R-CNN family, YOLOv3, and SSD. HRSC2016, DOTA Ships
Airplane Detection from VHR Images [13] Conducts a comprehensive evaluation of popular CNN-based models—YOLOv4, Faster R-CNN, and SSD—focusing on their performance in detecting airplanes from satellite imagery. The models were trained on the DOTA dataset and tested on both the DOTA and Pleiades satellite images. The study found that Faster R-CNN performed the best in terms of COCO metrics and F1 scores, excelling in object localization, particularly for larger airplanes. YOLOv4 demonstrated faster convergence and was more efficient in detecting smaller airplanes, making it suitable for real-time applications where processing speed is crucial. However, SSD, while less effective in detection performance, performed relatively well in object localization. DOTA, HRPlanes Airplanes
Ship Detection Using Wavelet Transforms [47] Focuses on improving ship detection and segmentation accuracy in optical remote sensing images by addressing common challenges such as background clutter, environmental noise, and false alarms. The authors propose an approach using wavelet transform to reduce the amount of false positives in the prediction, particularly in cluttered coastal areas. By incorporating multi-level false alarm identification, the method enhances the detection of ships within various challenging environments, such as cloud cover or water surface reflections. Custom dataset from coastal regions Ships
Small Objects in remote sensing [48] Focuses on the challenges and methods associated with detecting small objects in remote sensing imagery. It discusses the application of deep learning techniques to improve detection accuracy, especially for small-scale objects like vehicles, buildings, and other infrastructures in high-resolution satellite and aerial imagery. The review highlights the significance of using advanced architectures, such as YOLOv8 and Transformers. Various strategies are also subject of discussion such data augmentation and transfer learning. The authors also conducted a comparative study that includes 3 categories: one-stage, two-stage, and anchor-free. DOTA Multiclass: airplane, buildings, vehicles, etc.
@@ 3 Aircraft Datasets
In the field of aircraft detection and remote sensing, access to high-quality and diverse datasets is important for the development and evaluation of computer vision algorithms. In some comparative studies, the DOTA dataset [42] is selected, encompassing objects other than aircraft such as airports, bridges, and containers. This section reviews a collection of datasets specifically and exclusively designed for aircraft detection research, each offering unique features and advantages. An overview of these datasets is given in Table 3.
@@ 4 Literature Review: Object Detection
Many object detection models have been proposed in the past few decades, the majority of which are based on CNNs [55]. So far, researchers have classified these models into two main categories: one- and two-stage architectures [56, 57]. However, the recent rise of the self-attention mechanism [58] has led to a new category of network architectures based on Transformers. Figure 1 groups some of the most used and known models along with two types of classifications of deep learning models: (1) categorized based on their network architecture (one-stage, two-stage, or Transformer), or (2) based on their performance in real time and detection accuracy highlighted in blue and yellow, respectively. The overlapping area drawn in green features object detection models with balanced performances in both real-time detection and accuracy.
One-stage models [59] have been mainly known for their real-time deployment performance, because they are usually not computationally extensive and a single pass through the network is sufficient to produce estimations of object bounding boxes. However, the main limitation in this category is the detection accuracy, which might not be enough for some applications that require a very high detection confidence. In the case of two-stage object detection models, an additional stage is introduced to generate generic object proposals which make the model more efficient in its detection operation [60]. The purpose of this stage is to produce candidate bounding boxes, which may not be highly accurate, and effectively exclude background areas from further processing. Subsequently, the next stage of the model undertakes the more computationally intensive tasks of classifying objects and refining the bounding boxes generated by the previous stage. The third category is the Transformer-based architecture [61], which seems to be a good balance between the accuracy of real-time detection, as it has been tested and compared on common datasets. This category has been tested on common large datasets and it leverages the use of the self-attention mechanism to produce reliable and accurate bounding box estimations while being able to perform real-time detection.
The choice of object detection architecture depends on the application and the nature of images to process, which involves a trade-off between speed and accuracy [62]. One-stage models tend to offer faster processing speeds, but may exhibit lower accuracy compared to their two-stage counterparts. The advantages of one-stage algorithms can be summarized as follows [63, 64, 65]:
@@
  • Real-time Performance: One-stage detectors are designed to achieve real-time or near-real-time performance, making them suitable for applications where fast inference is crucial. These algorithms are commonly used in scenarios that require quick responses, such as autonomous driving, video analysis, and robotics [66].
  • Higher Recall: One-stage detectors tend to have a higher recall rate compared to two-stage detectors [67, 15]. They are capable of detecting a larger number of objects in an image, including small or occluded objects, because of their dense and dense-like prediction strategies. This higher recall can be advantageous in applications where comprehensive object detection is more important than achieving extremely precise bounding box localization.
@@
One-stageDetectionAccuracyReal-timeinferenceTransformerTwo-stages Overlap:
Real-time and accuracy
Deep learning based Object Detection RetinaNetCornerNetYOLOCenterNetRTMDetSSDRefineDetCasc.-RCNNR-FCNRCNNFast-RCNNFaster-RCNNMask-RCNNDETRRT-DETRViTDetGrounding DinoMViTv2
Figure 1: Classification of object detection methods based on (1) their architecture (one-stage, two-stage, and Transformer network), and (2) detection accuracy (in yellow) and real-time detection (blue). The red dot highlights the models that are implemented, trained, and validated in this work. The green outline (indicated as ‘Overlapping’ in the image) groups the models that usually perform well in both accuracy and inference time response.
@@
On the other hand, two-stage models generally achieve higher accuracy but sacrifice some speed due to the additional processing stage. The advantages of two-stage algorithms over the others can be summarized with the following points:
@@
  • Sampling Efficiency: Two-stage detectors employ a sampling mechanism to select a sparse set of region proposals, effectively eliminating a significant portion of the negative proposals. In contrast, one-stage detectors take a different approach by directly considering all regions in the image, which sometimes introduce class imbalance [68].
@@
  • Feature extraction: Two-stage detectors can allocate a larger head network for proposal classification and regression, as they only process a small number of proposals. This allows for the extraction of richer features, contributing to improved performance.
@@
  • Recall: Two-stage detectors take advantage of the RoIAlign [19, 69] operation to extract high-quality features from each proposal, ensuring location consistency. In contrast, in one-stage detectors, different region proposals may share the same feature, leading to coarse and spatially implicit representations that can cause feature misalignment.
@@
  • Accuracy: Two-stage detectors refine the location of the objects twice, once in each stage. Consequently, the bounding boxes generated by these models exhibit better accuracy compared to one-stage methods but at the expense of real-time performance.
@@
Consequently, the trade-off between one-stage, two-stage, and Transformer-based architectures necessitates careful consideration based on the specific requirements of the application at hand as well as the nature of images. In this context, special attention is directed towards the detection of airplanes within remote sensing images. The choice of algorithms is guided by their widespread usage, popularity in performance when tested on other types of images, and availability as open-source implementations. This will include YOLO, CenterNet, RTMDet, SSD, RetinaNet, Faster-RCNN, and DETR.
4.1 Models Backbones
Backbones are the foundational neural network architectures used for feature extraction in computer vision models [70], significantly influencing their accuracy and computational efficiency. They form the core of object detection and classification pipelines by providing hierarchical feature representations that are then used to identify objects both location and class at various scales and complexities. One of the first widely adopted backbones is the Very Deep Convolutional Network (VGG) [71], introduced by Simonyan and Zisserman, which uses deep convolutional layers to build robust feature maps. Despite its simplicity and effectiveness, VGG networks are computationally intensive, making them less suitable for real-time applications. To address this, ResNet [41] was proposed by He et al., introducing residual connections to ease deep network training and improve performance. ResNet variants, such as ResNet-50 and ResNet-101, remain fundamental in many object detection frameworks such as Faster R-CNN [18] and RetinaNet [33] due to their balance of efficiency and representational power. The field advanced later with feature pyramid designs, as seen in the Feature Pyramid Network (FPN) [72] layered atop backbones like ResNet. FPN enhances multi-scale feature extraction, making it particularly valuable for detecting objects of varying sizes, as implemented in YOLOv5 [73]. High-Resolution Networks (HRNet) [74] further improved upon this by maintaining high-resolution representations throughout the network, which proved advantageous for both semantic segmentation and object detection tasks. HRNet demonstrates high performance in models such as Mask R-CNN [19] and Cascade R-CNN [75].
Transformer-based architectures have emerged as another transformative innovation through Swin Transformers [76], which divide input images into patches and apply hierarchical attention mechanisms. These architectures are particularly effective in handling long-range dependencies and global context, making them prominent in modern frameworks like DETR [31] and DINO [77]. Swin Transformers also integrate well with dense prediction tasks, achieving state-of-the-art results across various benchmarks. Furthermore, YOLOv10 [28] represents one of the latest advances in the YOLO family for object detection, combining traditional convolutional approaches with Transformer-based modules to enhance attention mechanisms and reduce computational redundancy.
4.2 One-Stage Models
4.2.1 You Only Look Once